OpenStudy (anonymous):

Help please! Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4,1) (5,2) (7,2) (9,3)

OpenStudy (anonymous):

not understanding the whole question

OpenStudy (anonymous):

That's all there is to the problem

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

Okay

OpenStudy (dtan5457):

2x-5y=3 x-3y=1 All you need to do is solve this by elimination or substitution. I'll use elimination 2x-5y=3 x-3y=1 I'd multiply the 2nd equation by 2 2x-5y=3 2x-6y=2 Subtract and you get y=1 Now go solve for x by plugging in y=1 to any of your original equations.

OpenStudy (anonymous):

a is the answer

OpenStudy (anonymous):

Thank you!

OpenStudy (dtan5457):

@_tyannaj_ Once again, a full page explanation ruined by a direct answer that you clearly took my explanation. Yikes.